OpenStudy (anonymous):

find the area of an equilateral triangle with the give measurement 3 inch radius

OpenStudy (anonymous):

use the formula 1/2(bh)

OpenStudy (anonymous):

first find the hieght which is \[3\sqrt{3}\]

OpenStudy (anonymous):

the answer is 7.79 or \[9\sqrt{3}\]/2

OpenStudy (anonymous):

\[9\sqrt{3}/2\]

OpenStudy (anonymous):

Thanks